OpenStudy (anonymous):

how far apart are the points (3,-3) and (8,1) ?

OpenStudy (anonymous):

do you know the distance formula?

OpenStudy (anonymous):

i dont have one it doeesnt show it

OpenStudy (anonymous):

\[D=\sqrt{(x_2 - x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

all you need to do is subtract the x-coordinates then square that answer, do the same with the y-coordinates. square that anser . add them. then take the whole square root...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let's verify... what's your answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok.. subtract the x's. what are your x-coordinates?

OpenStudy (anonymous):

3-1= 2

OpenStudy (anonymous):

your x coordinates are 3 and 8...

OpenStudy (anonymous):

woops thats what i meant

OpenStudy (anonymous):

3-8=

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

good.. now subtract the y's

OpenStudy (anonymous):

-3-1=-4

OpenStudy (anonymous):

very nice... now watch..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

 can you tell me what goes in those boxes?